# Unified interpretation of 95 GeV di-photon and di-tau Excesses in the Georgi-Machacek Model

### v1

@@ Abstract
We revisit the 95 GeV excesses in the $\gamma\gamma$ and $\tau\tau$ channels in the Georgi–Machacek model, where a single light $CP$-even custodial singlet $H$ can account for both hints. Using one-loop renormalization group-improved effective potential and positive-definiteness conditions for vacuum stability, together with perturbative unitarity, electroweak precision tests, $B$-physics, and Higgs data, we identify a narrow but viable parameter region. In this region, charged and doubly charged scalars enhance $H\to\gamma\gamma$, while the fermionic rescaling controls $H\to\tau\tau$, yielding combined signal strengths compatible at the $2\sigma$ level. The scenario predicts characteristic patterns in the singlet mixing and triplet VEV and is highly testable at the HL-LHC and future lepton colliders via precision $\kappa_{V}$ measurements and direct exotic searches.
@@ 1 Introduction
The Georgi-Machacek (GM) model Georgi:1985nv ; Chanowitz:1985ug ; Gunion:1989ci ; Gunion:1990dt ; Gunion:1989we ; Haber:1999zh ; Godfrey:2010qb ; Logan:2010en ; Low:2012rj ; Killick:2013mya ; Efrati:2014uta ; Degrande:2015xnm extends the SM by introducing a complex ${\rm SU(2)}_{L}$ triplet scalar $\chi$ with hypercharge +1+1 and a real ${\rm SU(2)}_{L}$ triplet scalar $\xi$ with hypercharge 0. The field content is arranged such that custodial symmetry is preserved at tree level after EWSB, provided the triplet vacuum expectation values (VEVs) are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$. This alignment permits $v_{\Delta}$ values up to several tens GeV, significantly larger than in generic triplet extensions, while maintaining $\rho_{\rm tree}=1$. The model predicts nine additional scalar states, including a characteristic doubly charged Higgs boson, resulting in a rich phenomenology and distinctive collider signatures Chiang:2012cn ; Kanemura:2013mc ; Englert:2013zpa ; Englert:2013wga ; Godunov:2014waa ; Chiang:2015kka ; Chiang:2015amq ; Yu:2016skv ; Cao:2016rht ; Chang:2017niy ; Zhang:2017och ; Li:2017daq ; Sun:2017mue ; Azuelos:2017dqw ; Cao:2018xwy ; Yang:2018pzt ; Ghosh:2019qie ; Ismail:2020zoz ; Zhu:2020bew ; Wang:2022okq ; Ghosh:2022wbe ; deLima:2022yvn ; Ahriche:2022aoj ; Ghosh:2022bvz ; Chakraborti:2023mya ; Ghosh:2023izq ; Bhattacharya:2024gyh ; Lu:2024ade ; Ghosh:2025htt ; Ashanujjaman:2025puu ; Ghosh:2025hyy ; Chiang:2025lab ; Chiang:2018cgb . Notably, the GM model allows the coupling of the SM-like Higgs boson to weak gauge bosons to exceed the SM value Falkowski:2012vh ; Chang:2012gn ; Chiang:2013rua .
@@
The extended scalar sector improves vacuum stability and can induce strong first-order electroweak phase transitions (EWPT), potentially generating gravitational wave signals accessible to future experiments Chiang:2014hia ; Chen:2022zsh ; Zhou:2018zli ; Bian:2019bsn ; Lu:2025vif . If lepton number violation is present, triplet-lepton interactions can generate neutrino masses Chen:2020ark ; Giarnetti:2023dcr ; Giarnetti:2023osf ; Du:2022brr . The GM framework has also been employed to address the $W$-mass anomaly reported by CDF-II Du:2022brr ; Mondal:2022xdy ; Chen:2022ocr ; Ghosh:2022zqs , especially when custodial symmetry breaking effects are included Blasi:2017xmc ; Keeshan:2018ypw ; Song:2022jns . Furthermore, the light custodial singlet scalar of the GM model constitutes a natural candidate to interpret the excesses observed near 95 GeV Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop ; Ghosh:2025ebc ; Mondal:2025tzi . A unified explanation for the excesses observed in the $\gamma\gamma$, $\tau\tau$, and $b\bar{b}$ channels requires a near-degenerate CPCP-even/CPCP-odd ‘twin-peak’ structure Chen:2023bqr ; with only a single light $CP$-even state, the $\tau\tau$ excess is in tension with the $\gamma\gamma$ and $b\bar{b}$ channels Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop . Recent analyses, however, strongly disfavor the LEP $b\bar{b}$ excess as evidence for a new scalar Janot:2024ryq , motivating a focus on the LHC-driven $\gamma\gamma$ and $\tau\tau$ channels. In this context, the newly proposed positive-definiteness vacuum stability conditions based on one-loop renormalization group-improved (RG-improved) effective potential Du:2024mry expand the viable parameter space beyond that allowed by traditional bounded-from-below requirements. This enables a unified interpretation of both excesses within a single $CP$-even Higgs boson scenario.
@@
In this work, we focus on the unified interpretation of $\gamma\gamma$ and $\tau\tau$ resonance signals within the GM model. This work is organised as follows. In Sec. 2, we provide a brief review of the GM model. A short discussion about the di-photon and di-tau excess signals at a mass of 95 GeV is presented in Sec. 3. Theoretical and experimental constraints relevant to realistic research within the GM model are detailed in Sec. 4. Numerical results are displayed in Sec. 5. Finally, we present our conclusions in Sec. 6.
@@ 2 The Georgi-Machacek model
Therefore, when the triplet VEVs are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$, the vacuum remains invariant only under the diagonal (custodial) subgroup ${\rm SU(2)}_{V}\subset{\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ (i.e., when $U_{L}=U_{R}$). As a result, the tree-level $\rho$ parameter automatically satisfies $\rho_{\rm tree}=1$. In this alignment limit, the VEVs of scalar field take the form
@@
And the condition for EWSB is given as
@@
$\tan{2\alpha}=\frac{2M_{12}^{2}}{M_{22}^{2}-M_{11}^{2}},\quad\alpha\in\left(0,\pi\right)$ (16)
@@
$\begin{split}g_{hff}&=\kappa_{ff}^{h}\times g_{hff}^{\rm SM},\quad g_{hVV}=\kappa_{VV}^{h}\times g_{hVV}^{\rm SM},\\ g_{Hff}&=\kappa_{ff}^{H}\times g_{hff}^{\rm SM},\quad g_{HVV}=\kappa_{VV}^{h}\times g_{hVV}^{\rm SM},\end{split}$ (21)
@@
Using this framework, the GM prediction of the di-photon signal strength normalised to its SM prediction $\mu_{\gamma\gamma}$ are given by
$\begin{split}\mu_{\gamma\gamma}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\gamma\gamma\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\gamma\gamma\right)}\\ &\simeq\frac{\sigma_{\rm GM}\left(gg\to H_{95}\right)}{\sigma_{\rm SM}\left(gg\to h_{95}\right)}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{\mathcal{B}_{\rm SM}\left(h_{95}\to\gamma\gamma\right)}\\ &=\left|\kappa_{ff}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{1.39\times 10^{-3}},\end{split}$ (23)
where the mass of the Higgs boson is fixed around 95 GeV (denoted by $H_{95}$ and $h_{95}$), and the subscript $\rm GM$ ($\rm SM$) denotes the predictions of the GM model (SM) on the inclusive production rate of $H_{95}$($h_{95}$). The specific value $1.39\times 10^{-3}$ is the SM Higgs branching ratio at 95 GeV calculated by LHC Higgs Cross Section Working Group LHCHiggsCrossSectionWorkingGroup:2013rie . The second line approximation arises because the production of the SM Higgs boson is dominated by the gluon-gluon fusion process, which accounts for approximately 89% of the total Higgs production at a mass of 95 GeV LHCHiggsCrossSectionWorkingGroup:2013rie . The third line rises because the gluon-gluon fusion process ($gg\to h$) is mediated by a colored-fermion loop. As a result, the effective coupling $\kappa^{H}_{gg}$ equals $\kappa^{H}_{ff}$, since the fermionic couplings of $H$ scale universally as $\kappa_{ff}^{H}$ in lowest order, as shown in Eq. (22). Similarly, the GM prediction of the $\tau\tau$ and $bb$ channel signal strengths are given by
$\begin{split}\mu_{bb}&\equiv\frac{\sigma_{\rm GM}\left(e^{+}e^{-}\to ZH_{95}(H_{95}\to bb)\right)}{\sigma_{\rm SM}\left(e^{+}e^{-}\to Zh_{95}(h_{95}\to\gamma\gamma\right)}\\ &=\left|\kappa_{VV}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to bb\right)}{0.801},\\ \mu_{\tau\tau}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\tau\tau\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\tau\tau\right)}\\ &=\left|\kappa_{ff}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\tau\tau\right)}{8.32\times 10^{-2}}.\end{split}$ (24)
From above definitions, all strengths are expressed in term of the ratio $\sigma_{\rm GM}/\sigma_{\rm SM}$, so the QCD correction to the numerator and the denominator will cancel. This is beneficial to reduce the theoretical uncertainty in predicting the strengths.
@@
$\Gamma(H\to XX)=|\kappa^{H}_{XX}|^{2}\times\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}},$ (25)
@@
where $\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}}$ is the SM Higgs partial decay width estimated at mass of $m_{H}$. Therefore, the total decay width of custodial singlet state $H$ can be written as
$\Gamma^{\rm tot}_{H}=\Gamma^{\rm SM}_{h}\sum_{X}\left|\kappa^{H}_{XX}\right|^{2}\mathcal{B}_{\rm SM}\left(h\to XX\right),$ (26)
@@
where the SM Higgs total decay width $\Gamma^{\rm SM}_{h}$ and the branching rations $\mathcal{B}_{\rm SM}\left(h\to XX\right)$ estimated at $m_{H}$. The coupling between neutral scalar $H$ or $h$ and photons are induced by charged particle loops. In SM, the Higgs boson decay into two photons is mediated by $W$ boson and heavy charged fermion (top quark dominated) loops. While in GM model, the heavy charged Higgs states $S_{i}$ (including $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$ and $H_{3}^{\pm}$) loops can significantly enhance the partial decay width $\Gamma(H\to\gamma\gamma)$. The corresponding effective coupling $\kappa^{H}_{\gamma\gamma}$ can be expressed as
$\small\kappa^{H}_{\gamma\gamma}=\frac{\sum_{i}\dfrac{g_{HS_{i}S_{i}}}{2m_{S_{i}}^{2}}vQ_{S_{i}}^{2}A_{0}\left(\frac{m_{H}^{2}}{4m_{S_{i}}^{2}}\right)+\kappa_{VV}^{H}A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\kappa_{ff}^{H}~\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)}{A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)},$ (27)
@@
where $g_{HS_{i}S_{i}}$ are the scalar trilinear couplings of $H$ to the charged states $S_{i}$, and complex factors $A_{0,1/2,1}$ can be found in the literature Djouadi:2005gi ; Djouadi:2005gj ; Gunion:1989we . For an $H$ boson with mass $95~{\rm GeV}$, $A_{1}\simeq-7.6$ and $A_{1/2}\approx 4/3$. The factor $A_{0}$ reaches its maximum value of $1.5+i$ at $m_{H}=2m_{S_{i}}$, while in the limit of very heavy loop charged scalars, $A_{0}(0)\simeq 1/3$.
Then the signal strength in Eq. (23) and Eq. (24) can be rewriten as:
@@
$\begin{split}\mu_{\gamma\gamma}&=\left|\kappa^{H}_{ff}\right|^{2}\left|\kappa^{H}_{\gamma\gamma}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}\\ \mu_{bb}&=\left|\kappa^{H}_{VV}\right|^{2}\left|\kappa^{H}_{ff}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}},\\ \mu_{\tau\tau}&=\left|\kappa^{H}_{ff}\right|^{4}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}.\end{split}$ (28)
@@
It is clear that all three signal strengths share a common factor, $\left|\kappa^{H}_{ff}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}$. This introduces some tensions in attributing the observed excesses to a single source. While charged scalar states in GM models can provide a positive contributions to interpreting the $\gamma\gamma$ excess, especially the doubly charged $H_{5}^{\pm\pm}$ benefiting from $Q_{S}^{2}=4$. Ref. Chen:2023bqr investigate the different possibility of $H$ explaining $\gamma\gamma$, $\gamma\gamma+bb$ and $\gamma\gamma+bb+\tau\tau$. However, recent dedicated reanalyses argue that this feature is unlikely to originate from a genuine new scalar resonance Janot:2024ryq . Given the inferior dijet mass resolution at LEP and the updated conclusions, we do not attempt to accommodate the $b\bar{b}$ hint in this work. Instead, we focus on the two LHC-driven channels ($\gamma\gamma$ and $\tau\tau$) where the experimental systematics and resolutions are better controlled in the relevant mass window. Finally, we remind that all loop-induced rates, branching ratios and total widths in our scan are evaluated with the spectrum tools described in Sec. 4.3.

### v2

@@ Abstract
We revisit the 95 GeV excesses in the $\gamma\gamma$ and $\tau\tau$ channels in the Georgi–Machacek model, where a single light $CP$-even custodial singlet $H$ can account for both hints. Using one-loop renormalization group-improved effective potential and positive-definiteness conditions for vacuum stability, together with perturbative unitarity, electroweak precision tests, $B$-physics, and Higgs data, we identify a narrow but viable parameter region. In this region, charged and doubly charged scalars enhance $H\to\gamma\gamma$, while the fermionic rescaling controls $H\to\tau\tau$, yielding combined signal strengths compatible at the $2\sigma$ level. The scenario predicts characteristic patterns in the singlet mixing and triplet vacuum expectation values and is highly testable at the HL-LHC and future lepton colliders via precision $\kappa_{V}$ measurements and direct exotic searches.
@@ 1 Introduction
The Georgi-Machacek (GM) model Georgi:1985nv ; Chanowitz:1985ug ; Gunion:1989ci ; Gunion:1990dt ; Gunion:1989we ; Haber:1999zh ; Godfrey:2010qb ; Logan:2010en ; Low:2012rj ; Killick:2013mya ; Efrati:2014uta ; Degrande:2015xnm extends the SM by introducing a complex ${\rm SU(2)}_{L}$ triplet scalar $\chi$ with hypercharge +1+1 and a real ${\rm SU(2)}_{L}$ triplet scalar $\xi$ with hypercharge 0. The field content is arranged such that custodial symmetry is preserved at tree level after EWSB, provided the triplet vacuum expectation values (VEVs) are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$. This alignment permits $v_{\Delta}$ values up to several tens GeV, significantly larger than in generic triplet extensions, while maintaining $\rho_{\rm tree}=1$. The model predicts nine additional scalar states, including a characteristic doubly charged Higgs boson, resulting in a rich phenomenology and distinctive collider signatures Chiang:2012cn ; Kanemura:2013mc ; Englert:2013zpa ; Englert:2013wga ; Godunov:2014waa ; Chiang:2015kka ; Chiang:2015amq ; Yu:2016skv ; Cao:2016rht ; Chang:2017niy ; Zhang:2017och ; Li:2017daq ; Sun:2017mue ; Azuelos:2017dqw ; Cao:2018xwy ; Yang:2018pzt ; Ghosh:2019qie ; Ismail:2020zoz ; Zhu:2020bew ; Wang:2022okq ; Ghosh:2022wbe ; deLima:2022yvn ; Ahriche:2022aoj ; Ghosh:2022bvz ; Chakraborti:2023mya ; Ghosh:2023izq ; Bhattacharya:2024gyh ; Lu:2024ade ; Ghosh:2025htt ; Ashanujjaman:2025puu ; Ghosh:2025hyy ; Chiang:2025lab ; Chiang:2018cgb . Notably, the GM model allows the couplings of the SM-like Higgs boson to weak gauge bosons to exceed their SM values Falkowski:2012vh ; Chang:2012gn ; Chiang:2013rua .
@@
The extended scalar sector improves vacuum stability and can induce strong first-order electroweak phase transitions (EWPT), potentially generating gravitational wave signals accessible to future experiments Chiang:2014hia ; Chen:2022zsh ; Zhou:2018zli ; Bian:2019bsn ; Lu:2025vif . If lepton number violation is present, triplet-lepton interactions can generate neutrino masses Chen:2020ark ; Giarnetti:2023dcr ; Giarnetti:2023osf ; Du:2022brr . The GM framework has also been employed to address the $W$-mass anomaly reported by CDF-II Du:2022brr ; Mondal:2022xdy ; Chen:2022ocr ; Ghosh:2022zqs , especially when custodial symmetry breaking effects are included Blasi:2017xmc ; Keeshan:2018ypw ; Song:2022jns . Furthermore, the light custodial singlet scalar of the GM model constitutes a natural candidate to interpret the excesses observed near 95 GeV Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop ; Ghosh:2025ebc ; Mondal:2025tzi . A unified explanation for the excesses observed in the $\gamma\gamma$, $\tau\tau$, and $b\bar{b}$ channels requires a near-degenerate CPCP-even/CPCP-odd ‘twin-peak’ structure Ahriche:2023wkj ; with only a single light $CP$-even state, the $\tau\tau$ excess is in tension with the $\gamma\gamma$ and $b\bar{b}$ channels Ahriche:2023wkj ; Chen:2023bqr ; Du:2025eop . Recent analyses, however, strongly disfavor the LEP $b\bar{b}$ excess as evidence for a new scalar Janot:2024ryq , motivating a focus on the LHC-driven $\gamma\gamma$ and $\tau\tau$ channels.
The positive-definiteness vacuum stability conditions for the GM model were proposed in Du:2024mry , which preclude the existence of deeper vacua in large field value regions. Moving beyond the conventional tree-level analysis, the study employs the one-loop renormalization group-improved (RG-improved) effective potential, along with a novel criterion for the positive definiteness of homogeneous polynomials in multiple variables—a refinement necessitated by loop-induced custodial symmetry breaking effects. Through numerical analysis, it is shown that the quartic coupling ranges permitted by the positive-definiteness requirements significantly deviate from those derived from tree-level bounded-from-below conditions. Notably, certain parameter regions previously excluded by tree-level criteria become viable when the one-loop RG-improved scalar potential is considered. Conversely, some regions of parameter space that satisfy tree-level bounded-from-below constraints at the electroweak scale are ultimately ruled out by the positive-definiteness requirements of the effective potential in large field value regimes.
In this work, the newly proposed positive-definiteness vacuum stability conditions based on the one-loop RG-improved effective potential expand the viable parameter space beyond that allowed by traditional bounded-from-below requirements. This enables a unified interpretation of both excesses within a single $CP$-even Higgs-boson scenario, and we therefore focus on the unified interpretation of $\gamma\gamma$ and $\tau\tau$ resonance signals within the GM model.
@@
This work is organized as follows. In Sec. 2, we provide a brief review of the GM model. A short discussion of the di-photon and di-tau excess signals around 95 GeV is presented in Sec. 3. Theoretical and experimental constraints relevant to realistic studies within the GM model are detailed in Sec. 4. Numerical results are displayed in Sec. 5. Finally, we present our conclusions in Sec. 6.
@@ 2 The Georgi-Machacek model
Therefore, when the triplet VEVs are aligned, $v_{\chi}=v_{\xi}\equiv v_{\Delta}$, the vacuum remains invariant only under the diagonal (custodial) subgroup ${\rm SU(2)}_{V}\subset{\rm SU(2)}_{L}\times{\rm SU(2)}_{R}$ (i.e., when $U_{L}=U_{R}$). As a result, the tree-level $\rho$ parameter automatically satisfies $\rho_{\rm tree}=1$. In this alignment limit, the VEVs of the scalar fields take the form
@@
The condition for EWSB is given by
@@
$\tan{2\alpha}=\frac{2M_{12}^{2}}{M_{22}^{2}-M_{11}^{2}},\quad\alpha\in\left(-\pi/2,\pi/2\right)$ (16)
@@
$\begin{split}g_{hff}&=\kappa_{F}\times g_{hff}^{\rm SM},\quad g_{hVV}=\kappa_{V}\times g_{hVV}^{\rm SM},\\ g_{Hff}&=\kappa_{F}^{H}\times g_{hff}^{\rm SM},\quad g_{HVV}=\kappa_{V}^{H}\times g_{hVV}^{\rm SM},\end{split}$ (21)
@@
Using this framework, the GM predictions for the di-photon signal strength normalized to the SM prediction $\mu_{\gamma\gamma}$ are given by
$\begin{split}\mu_{\gamma\gamma}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\gamma\gamma\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\gamma\gamma\right)}\\ &\simeq\frac{\sigma_{\rm GM}\left(gg\to H_{95}\right)}{\sigma_{\rm SM}\left(gg\to h_{95}\right)}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{\mathcal{B}_{\rm SM}\left(h_{95}\to\gamma\gamma\right)}\\ &=\left|\kappa_{F}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\gamma\gamma\right)}{1.39\times 10^{-3}},\end{split}$ (23)
where the mass of the Higgs boson is fixed around 95 GeV (denoted by $H_{95}$ and $h_{95}$), and the subscript $\rm GM$ ($\rm SM$) denotes the GM (SM) prediction for the inclusive production rate of $H_{95}$ ($h_{95}$). The specific value $1.39\times 10^{-3}$ is the SM Higgs branching ratio at 95 GeV calculated by LHC Higgs Cross Section Working Group LHCHiggsCrossSectionWorkingGroup:2013rie . The second line approximation arises because the production of the SM Higgs boson is dominated by the gluon-gluon fusion process, which accounts for approximately 89% of the total Higgs production at a mass of 95 GeV LHCHiggsCrossSectionWorkingGroup:2013rie . The third line arises because the gluon-gluon fusion process ($gg\to h$) is mediated by a colored-fermion loop. As a result, the effective coupling $\kappa^{H}_{gg}$ equals $\kappa^{H}_{F}$, since the fermionic couplings of $H$ scale universally as $\kappa_{F}^{H}$ in lowest order, as shown in Eq. (22). Similarly, the GM predictions for the signal strengths in the $\tau\tau$ and $bb$ channels are given by
$\begin{split}\mu_{bb}&\equiv\frac{\sigma_{\rm GM}\left(e^{+}e^{-}\to ZH_{95}(H_{95}\to bb)\right)}{\sigma_{\rm SM}\left(e^{+}e^{-}\to Zh_{95}(h_{95}\to bb)\right)}\\ &=\left|\kappa_{V}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to bb\right)}{0.801},\\ \mu_{\tau\tau}&\equiv\frac{\sigma_{\rm GM}\left(pp\to H_{95}\to\tau\tau\right)}{\sigma_{\rm SM}\left(pp\to h_{95}\to\tau\tau\right)}\\ &=\left|\kappa_{F}^{H}\right|^{2}\times\frac{\mathcal{B}_{\rm GM}\left(H_{95}\to\tau\tau\right)}{8.32\times 10^{-2}}.\end{split}$ (24)
From the above definitions, all signal strengths are expressed in terms of the ratio $\sigma_{\rm GM}/\sigma_{\rm SM}$, so QCD corrections to the numerator and denominator largely cancel. This helps reduce the theoretical uncertainty in the predicted signal strengths.
@@
$\Gamma(H\to XX)=|\kappa^{H}_{X}|^{2}\times\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}},$ (25)
@@
where $\Gamma_{\rm SM}(h\to XX)\big|_{m_{h}=m_{H}}$ is the SM Higgs partial decay width evaluated at mass $m_{H}$. Therefore, the total decay width of custodial singlet state $H$ can be written as
$\Gamma^{\rm tot}_{H}=\Gamma^{\rm SM}_{h}\sum_{X}\left|\kappa^{H}_{X}\right|^{2}\mathcal{B}_{\rm SM}\left(h\to XX\right),$ (26)
@@
where the SM Higgs total decay width $\Gamma^{\rm SM}_{h}$ and branching ratios $\mathcal{B}_{\rm SM}\left(h\to XX\right)$ are evaluated at $m_{H}$. The couplings between the neutral scalars $H$ and $h$ and photons are induced by charged-particle loops. In the SM, Higgs-boson decay into two photons is mediated by $W$-boson and heavy charged-fermion (top-quark-dominated) loops. In the GM model, loops of heavy charged Higgs states $S_{i}$ (including $H_{5}^{\pm\pm}$, $H_{5}^{\pm}$, and $H_{3}^{\pm}$) can significantly enhance the partial decay width $\Gamma(H\to\gamma\gamma)$. The corresponding effective coupling $\kappa^{H}_{\gamma\gamma}$ can be expressed as
$\small\kappa^{H}_{\gamma}=\frac{\sum_{i}\dfrac{g_{HS_{i}S_{i}}}{2m_{S_{i}}^{2}}vQ_{S_{i}}^{2}A_{0}\left(\frac{m_{H}^{2}}{4m_{S_{i}}^{2}}\right)+\kappa_{V}^{H}A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\kappa_{F}^{H}~\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)}{A_{1}\left(\frac{m_{H}^{2}}{4m_{W}^{2}}\right)+\dfrac{4}{3}A_{1/2}\left(\frac{m_{H}^{2}}{4m_{t}^{2}}\right)},$ (27)
@@
where $g_{HS_{i}S_{i}}$ are the scalar trilinear couplings of $H$ to the charged states $S_{i}$, and the loop functions $A_{0,1/2,1}$ can be found in the literature Djouadi:2005gi ; Djouadi:2005gj ; Gunion:1989we . For an $H$ boson with mass $95~{\rm GeV}$, $A_{1}\simeq-7.6$ and $A_{1/2}\approx 4/3$. The factor $A_{0}$ reaches its maximum value of $1.5+\mathrm{i}$ at $m_{H}=2m_{S_{i}}$, while in the limit of very heavy loop charged scalars, $A_{0}(0)\simeq 1/3$.
Then the signal strength in Eq. (23) and Eq. (24) can be rewritten as:
@@
$\begin{split}\mu_{\gamma\gamma}&=\left|\kappa^{H}_{F}\right|^{2}\left|\kappa^{H}_{\gamma}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}\\ \mu_{bb}&=\left|\kappa^{H}_{V}\right|^{2}\left|\kappa^{H}_{F}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}},\\ \mu_{\tau\tau}&=\left|\kappa^{H}_{F}\right|^{4}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}.\end{split}$ (28)
@@
It is clear that all three signal strengths share a common factor, $\left|\kappa^{H}_{F}\right|^{2}\cdot\frac{\Gamma_{h}^{\rm SM}}{\Gamma^{\rm tot}_{H}}$. This introduces tension in attributing all observed excesses to a single source. Charged scalar states in the GM model can provide positive contributions to the $\gamma\gamma$ excess, especially the doubly charged $H_{5}^{\pm\pm}$ with $Q_{S}^{2}=4$. Ref. Chen:2023bqr investigates different possibilities for $H$ to explain $\gamma\gamma$, $\gamma\gamma+bb$, and $\gamma\gamma+bb+\tau\tau$. However, recent dedicated reanalyses argue that this feature is unlikely to originate from a genuine new scalar resonance Janot:2024ryq . Given the inferior di-jet mass resolution at LEP and the updated conclusions, we do not attempt to accommodate the $b\bar{b}$ hint in this work. Instead, we focus on the two LHC-driven channels ($\gamma\gamma$ and $\tau\tau$) where the experimental systematics and resolutions are better controlled in the relevant mass window. Finally, we note that all loop-induced rates, branching ratios and total widths in our scan are evaluated with the spectrum tools described in Sec. 4.